# thanks: 0009-0007-0442-5443thanks: 0000-0002-4641-022Xthanks: 0000-0001-9980-0363

### v2

@@ Abstract
Originally, the Biswas–Chatterjee–Sen model exhibits an order/disorder phase transition for a sufficiently large number of negative interactions among actors. In this paper, the model is extended by the nonconformists and inflexibles. Nonconformists are actors who do not follow the original model rules, but in different ways do something opposite. We introduce inflexibles as actors who does not change their opinions. Both discrete and continuous opinions are considered. With direct Monte Carlo simulations and mean-field calculations, we check the influence of fractions of nonconformists and inflexibles on mean opinion in the system. With the mean-field calculations we identify ranges of fractions of nonconformists where ordered phase of the system is available. The results of the mean-field calculations perfectly match the results of the Monte Carlo simulations. We consider inflexibles adhered: (i) to extreme opinions; (ii) to specific opinions and (iii) chosen independently of their initial opinion. For inflexibles adhered to specific and extreme opinions they play a role of effective bias suppressing disorder phase in the system. The qualitative results of introducing nonconformists (inflexibles) in various ways (discrete/continuous opinions and annealed/quenched disorder) are roughly the same. However, for the model extended by inflexibles, we can observe a systematic shift of the mean order parameter to its higher values for quenched disorder compared with annealed disorder.
@@ I Introduction
The opinion formation models may be classified into discrete or continuous ones. The latter are represented by the models of Weisbuch et al. [9, 10] or Hegselmann and Krause [11] applied also to complex networks [12, 13]. In such models, opinions are represented as real numbers, for example, taking values in the interval [1;+1][-1;+1]. Among discrete models one should mention: voter [14, 15, 16, 17], majority-rule [18], Sznajd [19] or models based on social impact [20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. In such models, the opinions are represented as integers, in the simplest case with only two opinions in $\{-1,+1\}$. However, for discrete models, also more than two opinion values (but still with their finite number) were investigated for the voter model [30, 31, 32, 33, 34, 35, 36, 29], the Sznajd model [37, 38, 39, 29], the Latané model [29, 28, 27, 25, 24], the majority-rule model [40, 41, 42, 43, 44] and in other studies [45, 46, 47, 48, 49].
The kinetic model of opinion dynamics proposed by Biswas, Chatterjee and Sen (BChS) [50, 51] was originally formulated in both versions: continuous and discrete. Soon after original introducing, the BChS model was followed by several modifications [52, 53] with the most recent presented in References 54, 55, 56, 57 (see also References 58, 59 for reviews).
@@
Using both mean-field theory calculations and direct Monte Carlo simulation results, we show how the fraction of nonconformists and the fraction of inflexibles influence the original model results. We consider both, discrete and continuous spectrum of opinions. Also various variants of disorder (annealed or quenched) are considered. For annealed disorder, every actor decides to behave as nonconformists (or inflexibles) with a given probability corresponding to the assumed fraction of them in the society. For quenched disorder, the positions of inflexibles and nonconformists are fixed at the network and their fraction are the model control parameters.
Some ways of introduction of inflexibles may shift the original curve of the order parameter vs. noise level towards higher values of the order parameter. As a result, the disordered phase is removed from the system.
In contrast, introducing nonconformists leads to the reduction of the order parameter. For some range of their abundance, only a disordered phase is observed. The results of performed mean-filed calculations perfectly match the results of direct Monte Carlo simulations.
@@ II Model
The order parameter of the model is defined as
@@ III.1 Mean-field calculations
III.2 Mean-field calculations for nonconformist (4)
We now introduce nonconformists in the first version according to Equation˜4 following the simplest rule: In an update, the nonconformist agents flips the sign of the influence, i.e., $\mu_{ij}\to-\mu_{ij}$ with probability $c$ and during a normal update with probability 1c1-c, the dynamics remains unchanged.
@@ III.2 Mean-field calculations for nonconformist (4)
In the original model, all mean-field transition rates depend on $\mu$ only through the probability that the influence is negative, namely $p$. In the presence of flip-$\mu$ nonconformists, a fraction $c$ of the updates uses the opposite sign, so the probability that $\mu$ is negative becomes $p_{\text{eff}}$. Therefore, the mean-field evolution in the presence of nonconformists is obtained exactly by replacing $p\to p_{\text{eff}}$. This observation allows us to use the known mean-field equations of the original model while incorporating nonconformists through the single parameter $p_{\text{eff}}$.
@@
where $p_{\text{eff}}$ can be found from Equation˜9. Hence ordered state exists if $p_{\text{eff}}<1/4$.
@@ III.2.1 Stability of the disordered fixed point
The first eigenvalue is always negative, ensuring stability along the $s$ direction. The second eigenvalue changes sign at $p_{\text{eff}}=1/4$. For $p_{\text{eff}}>1/4$, both eigenvalues are negative and the disordered state is linearly stable, whereas for $p_{\text{eff}}<1/4$, the eigenvalue associated with the order parameter direction becomes positive, making the point a saddle. Thus the disordered fixed point loses stability precisely at $p_{\text{eff}}=1/4$, which identifying the critical line separating the ordered and disordered phases.
@@
For $p<p_{c}(c)$, the condition $p_{\text{eff}}<1/4$ is reached and the disordered fixed point becomes a saddle. At this point a pair of ordered fixed points with nonzero $\pm O^{*}$ [Equation˜15] emerges and becomes stable. Thus, in the interval $0<c<1/4$, the system displays a genuine order-disorder transition at $p=p_{c}(c)$, with ordered state present whenever $p$ is sufficiently small.
@@ III.2.3 Divergence of the relaxation timescale
III.3 Mean-field calculations for nonconformists (5)
@@ III.5.2 Continuous opinions
Figure 5: Mean order parameter $\langle\{O\}\rangle$ for various fractions $c$ of nonconformists with discrete opinions modeled according to Equation˜4. The solid line represent the closed form expression of the order parameter given in Equation˜15 while the points denote the simulation results. The analytical prediction shows excellent agreement with the simulation data. All simulations were done for the system size $N=1024$
Figure 6: Mean order parameter $\langle\{O\}\rangle$ for various fractions $c$ of nonconformists with discrete opinions modeled according to Equation˜5. The solid line represent the closed form expression of the order parameter given in Equation˜41 while the points denote the simulation results. The analytical prediction shows excellent agreement with the simulation data. All simulations were done for the system size $N=1024$
@@ IV Discussion
For nonconformists introduced according to Equation˜4 the mean-field approach leads to conclusions that an ordered phase is available when the fraction of nonconformists is $c<1/4$ or $c>3/4$. The mean-field calculations applied to the nonconformists introduced according to Equation˜5 show that the ordered state may be reached only when $c<1/8$.
@@
For inflexibles adhered to extreme opinions (Section˜III.4.1), the symmetric point $\rho=1$ (with the fractions of inflexibles adhered to opinion +1+1 equal to the fractions of inflexibles adhered to opinion 1-1) is special. It is the only value of $\rho$ for which the order parameter equation is invariant under $O\to-O$ and admits a disordered fixed point. Any infinitesimal deviation from $\rho=1$ destroys this symmetry and suppresses the phase transition. In this sense, $\rho=1$ is a structurally unstable point protected by symmetry that separates unbiased dynamics from biased ones.
The introduction of inflexibles (Section˜III.4.2) adhered to a specific opinion (for example, +1+1) leads to a subsequent increase in the order parameter with the fraction $z$ of them (up-to $O\equiv 1$, independently on $p$, for $z=1$). In such a case, the order disorder phase transition is suppressed (see Figure˜4).
For inflexibles chosen independently of their initial opinion (Section˜III.4.3) the mean order parameter $\langle\{O\}\rangle$ is independent of their fraction $z$ as predicted by mean-field calculations (84) as presented in Figure˜7.
@@
Figure˜8 shows the mean order parameter $\langle\{O\}\rangle$ for fractions $z=0.2$ (Figure˜8, left panel), $z=0.5$ (Figure˜8, middle panel) and $z=0.8$ (Figure˜8, right panel) of inflexibles and discrete/continuous opinions and the quenched/annealed disorder. For continuous opinions, only quenched disorder is possible to simulate. Independently of the fraction $z$ of inflexibles, the variant with quenched disorder outperforms the variant with annealed disorder to reach unanimity of opinions (i.e., increasing $\langle\{O\}\rangle$). In version with annealed disorder, every actor my suddenly (with probability 1z1-z) leave his/her opinion (for example +1+1), while for quenched disorder (with fixed positions of inflexibles) inflexibles extreme opinions never change, and thus reaching higher values of $\langle\{O\}\rangle$ is easier.

### v3

@@ Abstract
Originally, the Biswas–Chatterjee–Sen model was shown to exhibit an order/disorder phase transition for a sufficiently large number of negative interactions among actors. In this paper, the model is extended by the existence of nonconformists and inflexible. Nonconformists are actors who do not follow the original model rules, but in different ways do something opposite while inflexibles are those who do not change their opinions. Both discrete and continuous opinions are considered. With direct Monte Carlo simulations and mean-field calculations, we check the influence of fractions of nonconformists and inflexibles on the mean opinion in the system. With the mean-field calculations we identify ranges of fractions of nonconformists where ordered phase of the system is available. The results of the mean-field calculations perfectly match the results of the Monte Carlo simulations. We consider inflexibles adhered: (i) to extreme opinions; (ii) to specific opinions and (iii) chosen independently of their initial opinion. For inflexibles adhered to specific and extreme opinions they play a role of effective bias suppressing disorder phase in the system. The qualitative results of introducing nonconformists (inflexibles) in various ways (discrete/continuous opinions and annealed/quenched disorder) are roughly the same. However, for the model extended by inflexibles, we can observe a systematic shift of the mean order parameter to its higher values for quenched disorder compared with annealed disorder.
@@ I Introduction
The opinion formation models may be classified into discrete or continuous ones. The latter are represented by the models of Weisbuch et al. [9, 10] or Hegselmann and Krause [11], applied also to complex networks [12, 13]. In such models, opinions are represented as real numbers, for example, taking values in the interval [1;+1][-1;+1]. Among discrete models one should mention: voter [14, 15, 16, 17], majority-rule [18], Sznajd [19] or models based on social impact [20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. In such models, the opinions are represented as integers, in the simplest case with only two opinions in $\{-1,+1\}$. However, for discrete models, also more than two opinion values (but still with their finite number) were investigated for the voter model [30, 31, 32, 33, 34, 35, 36, 29], the Sznajd model [37, 38, 39, 29], the Latané model [29, 28, 27, 25, 24], the majority-rule model [40, 41, 42, 43, 44] and in other studies [45, 46, 47, 48, 49].
The kinetic model of opinion dynamics proposed by Biswas, Chatterjee and Sen (BChS) [50, 51] was originally formulated in both versions: continuous and discrete. Soon after its introduction, the BChS model was followed by several modifications [52, 53] with the most recent presented in References 54, 55, 56, 57 (see also References 58, 59 for reviews).
@@
Using both mean-field theory calculations and direct Monte Carlo simulation results, we show how the fraction of nonconformists and the fraction of inflexibles separately influence the original model results. We consider both, discrete and continuous spectrum of opinions. Also various variants of disorder (annealed or quenched) are considered. For annealed disorder, every actor decides to behave as nonconformists (or inflexibles) with a given probability corresponding to the assumed fraction of them in the society. For quenched disorder, the positions of inflexibles and nonconformists are fixed at the network and their fractions are the model control parameters.
Some ways of introduction of the inflexibles may shift the original curve of the order parameter versus noise level towards higher values of the order parameter. As a result, the disordered phase is removed from the system.
In contrast, introducing nonconformists leads to the reduction of the order parameter. For some range of their abundance, only a disordered phase is observed. The results of the mean-filed calculations perfectly match the results of direct Monte Carlo simulations.
@@ II Model
The order parameter of the model with $N$ actors in total is defined as
@@ III.1 Mean-field calculations
III.2 Mean-field calculations for nonconformist for the dynamics given by Equation˜4
We now introduce nonconformists in the first version according to Equation˜4 following the simplest rule: In an update, the nonconformist agents flips the sign of the influence, i.e., $\mu_{ij}\to-\mu_{ij}$ with probability $c$ and during a normal update with probability 1c1-c, the dynamics remain unchanged.
@@ III.2 Mean-field calculations for nonconformist for the dynamics given by Equation˜4
In the original model, all mean-field transition rates depend on $\mu$ only through the probability that the influence is negative, namely $p$. In the presence of flip-$\mu$ nonconformists, a fraction $c$ of the updates uses the opposite sign, so the effective probability that $\mu$ is negative becomes $p_{\text{eff}}$. Therefore, the mean-field evolution in the presence of nonconformists is obtained exactly by replacing $p\to p_{\text{eff}}$. This observation allows us to use the known mean-field equations of the original model while incorporating nonconformists through the single parameter $p_{\text{eff}}$.
@@
where $p_{\text{eff}}$ can be found from Equation˜9. Hence the ordered state exists if $p_{\text{eff}}<1/4$.
@@ III.2.1 Stability of the disordered fixed point
The first eigenvalue is always negative, ensuring stability along the $s$ direction. The second eigenvalue changes sign at $p_{\text{eff}}=1/4$. For $p_{\text{eff}}>1/4$, both eigenvalues are negative and the disordered state is linearly stable, whereas for $p_{\text{eff}}<1/4$, the eigenvalue associated with the order parameter direction becomes positive, making the point a saddle. Thus the disordered fixed point loses stability precisely at $p_{\text{eff}}=1/4$, which can be identified as the critical point separating the ordered and disordered phases.
@@
For $p<p_{c}(c)$, the condition $p_{\text{eff}}<1/4$ is reached and the disordered fixed point becomes a saddle. At this point a pair of ordered fixed points with nonzero $\pm O^{*}$ [Equation˜15] emerge and become stable. Thus, in the interval $0<c<1/4$, the system displays a genuine order-disorder transition at $p=p_{c}(c)$, with the ordered state present whenever $p$ is sufficiently small.
@@ III.2.3 Divergence of the relaxation timescale
III.3 Mean-field calculations for nonconformists for the dynamics given by Equation˜5
@@ III.5.2 Continuous opinions
Figure 5: Mean order parameter $\langle\{O\}\rangle$ for various fractions $c$ of nonconformists with discrete opinions modeled according to Equation˜4. The solid lines represent the closed form expression of the order parameter given in Equation˜15 while the points denote the simulation results. The analytical prediction shows excellent agreement with the simulation data. All simulations were done for the system size $N=1024$
Figure 6: Mean order parameter $\langle\{O\}\rangle$ for various fractions $c$ of nonconformists with discrete opinions modeled according to Equation˜5. The solid lines represent the closed form expression of the order parameter given in Equation˜41 while the points denote the simulation results. The analytical prediction shows excellent agreement with the simulation data. All simulations were done for the system size $N=1024$
@@ IV Discussion
Figure 7: Mean order parameter $\langle\{O\}\rangle$ for various probabilities $z_{+}$ and $z_{-}$ of inflexibles adhered to extreme opinions (+1+1 and 1-1, respectively). The upper panel corresponds to $\rho=1$ ($z_{+}=z_{-}$, Case I in Section˜III.4.1) while the bottom panel shows results for $\rho>1$ ($z_{+}>z_{-}$, Case II in Section˜III.4.1). Discrete opinions and annealed disorder are assumed. The solid line represents the closed form expression of the order parameter given in Equation˜84 while the points denote the simulation results. All simulations were done for the system size $N=1024$
@@
For nonconformists introduced according to Equation˜4 the mean-field approach leads to conclusions that an ordered phase is available when the fraction of nonconformists is $c<1/4$ or $c>3/4$. This is well visible in Figure˜1 (and also in Figure˜5, where solid lines represent the closed form expression of the order parameter given in Equation˜15). At the same time, both figures reveal that replacing $c\to(1-c)$ with a simultaneous replacement $p\to(1-p)$ leaves the shapes of the mean order parameter $\langle\{O\}\rangle$ vs. $p$ unchanged. This is directly related to the symmetry of Equation˜18 with respect to such a change of variables.
The mean-field calculations applied to the nonconformists introduced according to Equation˜5 show that the ordered state may be reached only when $c<1/8$.
@@
For inflexibles adhered to extreme opinions (Section˜III.4.1), the symmetric point $\rho=1$ (with the fractions of inflexibles adhered to opinion +1+1 equal to the fractions of inflexibles adhered to opinion 1-1, see LABEL:{fig:inflexibles_o_vs_p_annealed_cases12} (upper panel)) is special. It is the only value of $\rho$ for which the order parameter equation is invariant under $O\to-O$ and admits a disordered fixed point. Any infinitesimal deviation from $\rho=1$ (as presented in the bottom panel of LABEL:{fig:inflexibles_o_vs_p_annealed_cases12} for $z_{+}>z_{-}$) destroys this symmetry and suppresses the phase transition. In this sense, $\rho=1$ is a structurally unstable point protected by symmetry that separates unbiased dynamics from biased ones.
The introduction of inflexibles (Section˜III.4.2) adhered to a specific opinion (for example, +1+1) leads to a subsequent increase in the order parameter with the fraction $z$ of them (up-to $O\equiv 1$, independently on $p$, for $z=1$). In such a case, the order disorder phase transition is suppressed (see Figure˜2).
For inflexibles chosen independently of their initial opinion (Section˜III.4.3) the mean order parameter $\langle\{O\}\rangle$ is independent of their fraction $z$ as predicted by mean-field calculations (84) as presented in Figure˜8.
@@
Figure˜9 shows the mean order parameter $\langle\{O\}\rangle$ for fractions $z=0.2$ (Figure˜9, left panel), $z=0.5$ (Figure˜9, middle panel) and $z=0.8$ (Figure˜9, right panel) of inflexibles and discrete/continuous opinions and the quenched/annealed disorder. For continuous opinions, only quenched disorder is possible to simulate. Independently of the fraction $z$ of inflexibles, the variant with quenched disorder outperforms the variant with annealed disorder to reach unanimity of opinions (i.e., increasing $\langle\{O\}\rangle$). In version with annealed disorder, every actor my suddenly (with probability 1z1-z) leave his/her opinion (for example +1+1), while for quenched disorder (with fixed positions of inflexibles) inflexibles extreme opinions never change, and thus reaching higher values of $\langle\{O\}\rangle$ is easier.